OpenStudy (anonymous):

A parallelogram has vertices (-1,2), (4,4), (2,-1) and (-3,-3). What is the parallelogram and what is the area and perimeter of it?

OpenStudy (mokeira):

plot it

OpenStudy (anonymous):

When I plotted it was a rhombus...now all I have to find out is the area and perimeter of it so can you help me with that ?

OpenStudy (mokeira):

yea sure...let me try and explain

OpenStudy (mokeira):

since it is a rhombus, all sides are equal. Step 1: Find the length of a side \[\sqrt{(x _{1}-x _{2})^{2}+(y _{1}-y _{2})^{2}}\] \[\sqrt{(4--1)^{2}+(4-1)^{2}}\] \[\sqrt{(25)+(4)}\] \[\sqrt{29}\] Area= sidexside \[\sqrt{29}\times \sqrt{29}=29\] perimeter=\[\sqrt{29}+\sqrt{29}+\sqrt{29}+\sqrt{29}=4\sqrt{29}\]

OpenStudy (mokeira):

hope you have understood

OpenStudy (anonymous):

But isn't there like a specific equation for a rhombus or is that the equation just like the area and perimeter of a square?

OpenStudy (mokeira):

oh yeah..sorry!

OpenStudy (mokeira):

you have the length of the sides, so i guess you can get the diagonals multiply the diagonals then divide by 2